### v1

# NVS-Solver: Video Diffusion Model as Zero-Shot Novel View Synthesizer

Abstract
By harnessing the potent generative capabilities of pre-trained large video diffusion models, we propose NVS-Solver, a new novel view synthesis (NVS) paradigm that operates without the need for training. NVS-Solver adaptively modulates the diffusion sampling process with the given views to enable the creation of remarkable visual experiences from single or multiple views of static scenes or monocular videos of dynamic scenes. Specifically, built upon our theoretical modeling, we iteratively modulate the score function with the given scene priors represented with warped input views to control the video diffusion process. Moreover, by theoretically exploring the boundary of the estimation error, we achieve the modulation in an adaptive fashion according to the view pose and the number of diffusion steps. Extensive evaluations on both static and dynamic scenes substantiate the significant superiority of our NVS-Solver over state-of-the-art methods both quantitatively and qualitatively. Source code in https://github.com/ZHU-Zhiyu/NVS__\__Solver.
1 Introduction
In the realm of computer vision and graphics, novel view synthesis (NVS) from limited visual data remains a formidable challenge with profound implications across various domains, from entertainment [46, 16] to autonomous navigation [1, 24] and beyond [3, 62, 35, 30, 20]. However, addressing this challenge demands a sophisticated method capable of extracting meaningful information from sparse visual inputs and synthesizing coherent representations of unseen viewpoints [63, 58]. In this context, the emerging field of deep learning has witnessed remarkable strides, particularly with the advent of advanced generative models [48, 7, 11].
Recently, diffusion models [13, 41, 45] have garnered significant attention due to their exceptional ability to synthesize visual data. A prominent area of focus within this domain is video diffusion models [5, 21, 14, 31, 18, 21, 53], which have gained popularity for their remarkable video generation capabilities.
In this paper, we address the problem of novel view synthesis from single or multiple views of static scenes or monocular videos of dynamic scenes, leveraging the pre-trained large video diffusion model without additional training. Specifically, we theoretically formulate the NVS-oriented diffusion process as guided sampling, in which the intermediate diffusion results are modulated with the scene information from the given views. Moreover, we empirically and theoretically investigate the potential distribution of the error map to achieve adaptive modulation in the reverse diffusion process, with a reduced estimation error boundary.
In summary, the main contributions of this paper lie in:
  • we propose a new training-free novel view synthesis paradigm by leveraging pre-trained video diffusion models;
  • we theoretically formulate the process of adaptively utilizing the given scene information to control the video diffusion process; and
  • we demonstrate the remarkable performance of our paradigm under various scenarios.
2 Related Work
Diffusion Model indicates a kind of deep generative model [39], which is inspired by non-equilibrium statistical physics by iteratively appending noise into image data and then reversely removing noise and transferring to noise-free data distribution. Ho et al. [13] proposed a variance preserving (VP) diffusion denoising probabilistic model via progressively removing noise. Song et al. [45, 43, 44, 42] proposed a score-based image generation model by iteratively calculating the derivative of data distribution and utilizing the stochastic differential equation (SDE) [2] or ordinary differential equation (ODE)-based solvers [27, 45] to reverse the noise-adding process and derive the clean image distribution. Inspired by the success of image diffusion models, many works attempted to build video diffusion model to directly achieve video generation prompted by text [21, 53] or a single image [5, 18].
Diffusion sampling algorithm aims to speed up or control the diffusion process via regularizing or re-directing the reverse trajectory of diffusion models [25, 26, 61, 60, 50, 6, 9]. Song et al. [41] proposed denoising diffusion implicit models (DDIM) to accelerate the diffusion sampling by jumping to the clean image-space at each step. Lu et al. [25, 26, 61] proposed a series of fast ODE diffusion solvers given an analytic solution of ODE by its semi-linear nature. Moreover, the integration of non-linear network-related parts was further approximated by its Taylor series. Zhang et al. [60] explored the huge variance of distribution shift and then decoupled an exponential variance component from the score estimation model, thus reducing the discretization error. Chung et al. [9] proposed to regularize the intermediate derivative from the reconstruction process to achieve image restoration. Wang et al. [50] decoupled the image restoration into range-null spaces and focus on the reconstruction of null space, which contains the degraded information.
Novel view synthesis (NVS) targets at generating images of a scene from viewpoints not presented in the original data and has been a subject of extensive research in computer vision and graphics [33, 8, 47, 3, 36]. Early methods in this domain often relied on geometric approaches, such as multi-view stereo reconstruction [38, 17] and structure-from-motion [37, 32, 32] techniques, to synthesize novel viewpoints from multiple images captured from different angles. The advent of deep learning has revolutionized the field of NVS, enabling the synthesis of realistic images from pre-trained feature volumes. Neural Radiance Fields (NeRF) [30, 34, 4, 28, 23] enabled stunningly detailed renders from 2D images through volumetric rendering. Additionally, differentiable rendering has allowed gradients of rendering outputs with respect to scene parameters, facilitating direct optimization of scene geometries, lighting, and materials. Recently, 3D Gaussian Splatting [20] presented an explicit representation of the scene.
While these methods have shown promising results, they often suffer from limitations such as dependence on dense scene geometry, challenges in handling complex scene dynamics, and being hard to generalize [20]. Moreover, they require significant computational resources and suffer from artifacts such as disocclusions and view-dependent effects.
In light of these challenges, our work builds upon the strengths of recent advancements in deep learning-based NVS, with a focus on leveraging the robust zero-shot view synthesis capabilities of the video diffusion model. By harnessing latent representations derived from sparse or single-view inputs, our approach aims to overcome the limitations of existing methods and produce high-quality novel views with improved realism.
3 Preliminary
We briefly introduce some preliminary knowledge of diffusion models [39], which facilitates our subsequent analyses. We also refer readers to [43, 45] for more details. Generally, the forward SDE process of the latent diffusion model can be formulated as
$d\bm{x}=f(t)\bm{x}dt+g(t)d\bm{w},$ (1)
where $\bm{x}$ is the noised latent state, $t$ for the timestamp of diffusion, and the two scalar functions $f(t)$ and $g(t)$ output drift and diffusion coefficients, indicating the variation of data and noise components during the diffusion process, respectively. Naturally, we have the following ODE solution:
$d\bm{x}=\left[f(t)\bm{x}-\frac{1}{2}g^{2}(\bm{x})\nabla_{\bm{x}}\log(q_{t}(\bm{x}))\right]dt.$ (2)
Through training a score model $S_{\bm{\theta}}(\bm{x},t)$ parameterized with $\bm{\theta}$ to approximate the $\nabla_{\bm{x}}\log[q(\bm{x})]$ as
$\mathcal{L}=\gamma(t)\|S_{\bm{\theta}}(\bm{x},t)-\nabla_{\bm{x}}\log[q_{t}(\bm{x})]\|_{2}^{2},$ (3)
with $\gamma(t)>0$, we can calculate the clean image $\bm{x}_{0}$ via utilizing the score function $S_{\bm{\theta}}(\bm{x},t)$ to calculate the data distribution gradient as
$d\bm{x}=\left[f(t)\bm{x}-\frac{1}{2}g^{2}(\bm{x})S_{\bm{\theta}}(\bm{x},t)\right]dt.$ (4)
Moreover, since the noise of the diffusion process is usually parameterized by i.i.d. Gaussian noises $\mathcal{N}(\bm{\epsilon}_{t};\bm{0},\sigma(t)\mathbf{I})$ with a variance of $\sigma(t)$, the above diffusion process can be calculated as
$d\bm{x}=\left[f(t)\bm{x}-\frac{1}{2}g^{2}(\bm{x})\frac{\bm{\mu}_{t}-\bm{x}}{\sigma^{2}(t)}\right]dt,$ (5)
where $\bm{\mu}_{t}=\mathcal{X}_{\bm{\theta}}(\bm{x}_{t},t)$ is the estimated clean image from the noised image $\bm{x}_{t}$ at step $t$ by the denosing U-Net $\mathcal{X}_{\bm{\theta}}(\cdot,~{}\cdot)$ [^1] For simplicity, here we combine some parameterizations stemmed from EDM [19] into $\mathcal{X}_{\bm{\theta}}(\cdot)$.. Finally, considering the special case of variance exploding (VE) diffusion process [45] of the stable video diffusion (SVD) [5], Eq. (5) can be further simplified as
$d\bm{x}=\frac{\bm{x}-\bm{\mu}_{t}}{\sigma(t)}d\sigma(t).$ (6)
4 Proposed Method
Motivated by the powerful generative capability with realistic and consistent frames by large video diffusion models, we aim to adapt the pre-trained video diffusion model to the task of NVS without any additional training, leading to NVS-Solver. Generally, NVS-Solver focuses on effectively incorporating the prior information of the given scene to guide the reverse sampling of a video diffusion process towards high-quality and desired novel views. Specifically, we first theoretically reformulate NVS-oriented reverse diffusion sampling by modulating the score function with the given views (Sec. 4.1). Based on theoretically exploring the boundary of the diffusion estimation and depth-based warping errors, we propose to adaptively modulate the prior of the given view into diffusion process to reduce the potential boundary of the estimation error (Sec. 4.2).
In the following, we take the single view-based NVS to illustrate our method, as summarized in Algorithm 1, which involves from a given view $\mathbf{X}_{0,\mathbf{p}_{0}}$ with pose $\mathbf{p}_{0}$, reconstructing $N-1$ novel views at target poses $\{\mathbf{p}_{1},\cdots,\mathbf{p}_{i},\cdots,\mathbf{p}_{N-1}\}$, denoted as $\mathbf{X}_{0,\mathbf{p}_{i}}$. Note that our method is suitable for scenarios involving NVS from multiple views, as well as from monocular videos, as explained in Sec. 4.3. Also, we consider the pre-trained image-to-video diffusion model SVD [5].
Notations. Let $\mathbf{X}_{t}\in\mathbb{R}^{H\times W\times N}$ be the spatial-temporal latent of a set of $N$ images/views at the $t$-th diffusion step, $\mathbf{p}_{i}\in\mathbb{R}^{5}$ the pose of the $i$-th image ($10\leq i\leq N-$ 1), $\mathcal{X}_{\bm{\theta}}(\mathbf{X}_{t},t)$ a U-Net denoising $\mathbf{X}_{t}$ to estimate the clean image set $\bm{\mu}_{t}\in\mathbb{R}^{H\times W\times N}$, $\mathbf{X}_{t,\mathbf{p}_{i}}\in\mathbb{R}^{H\times W}$ a typical latent at step $t$, time $i$, and spatial pose $\mathbf{p}_{i}$, and $\bm{\mu}_{t,\mathbf{p}_{i}}=\mathcal{X}^{\mathbf{p}_{i}}_{\bm{\theta}}(\mathbf{X}_{t},t)$, indicating a intermediate estimation of clean image $\mathbf{X}_{0,\mathbf{p}_{i}}$.
4.1 Scene Prior Modulated Reverse Diffusion Sampling
Based on the formulation of the image reverse diffusion sampling process in Eq. (6), we have
$d\mathbf{X}_{t,\mathbf{p}_{i}}=\left[\frac{\mathbf{X}_{t,\mathbf{p}_{i}}-\mathcal{X}_{\bm{\theta}}^{\mathbf{p}_{i}}(\mathbf{X}_{t},t)}{\sigma(t)}\right]d\sigma(t).$ (7)
According to the intensity function [29], we can formulate the relationship between $\mathbf{X}_{0,\mathbf{p}_{0}}$ and $\mathbf{X}_{0,\mathbf{p}_{i}}$ through Taylor expansion:
$\mathbf{X}_{0,\mathbf{p}_{i}}=\mathcal{I}(\mathbf{p}_{0})+\frac{d\mathcal{I}(\mathbf{p})}{d\mathbf{p}}\Delta\mathbf{p}+\mathcal{O}^{2}(\Delta\mathbf{p}),$ (8)
where $\mathcal{I}(\cdot)$ is the intensity function, $\Delta\mathbf{p}:=\mathbf{p}_{i}-\mathbf{p}_{0}$ is the pose variation, and $\mathcal{O}^{2}(\Delta\mathbf{p})$ is the high order Taylor expansions. Based on the depth-driven image warping operation, we further have
Algorithm 1 NVS-Solver for Synthesizing Novel Views from Single Images
1:Input: given view $\{\mathbf{X}_{\mathbf{p}_{0}}\}$, estimated depth map $\widehat{\mathbf{D}}_{\mathbf{p}_{0}}$, target view poses $\{\mathbf{p}_{0},\cdots,\mathbf{p}_{N-1}\}$, and diffusion U-Net $\mathcal{X}_{\bm{\theta}}(\cdot)$.
2:Derive the warping component $\{\widehat{\mathbf{X}}_{0,\mathbf{p}_{0}},\cdots,\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}},\cdots,\widehat{\mathbf{X}}_{0,\mathbf{p}_{N-1}}\}$, using the given views, corresponding poses and depth maps by Eq. (10). $\triangleright$ Preparing diffusion guidance images.
3:Initialize $\mathbf{X}_{T}\sim\mathcal{N}(\mathbf{0},\sigma_{t}\mathbf{I})$
4:For $t=T,...,1$ do $\triangleright$ Video diffusion reverse sampling steps.
5:      Diffusion network forward $\bm{\mu}_{t}=\mathcal{X}_{\bm{\theta}}(\mathbf{X}_{t},t)$.
6:      For $i=0,\cdots,N-1$ do
7:            Calculate weights $\widehat{\lambda}(t,\mathbf{p})$ via Eq. (17) and $\widetilde{\bm{\mu}}_{t,\mathbf{p}}$ via Eq. (12).
8:            If Directly guided sampling then
9:                  Reverse $\mathbf{X}_{t,\mathbf{p}}$ to $\mathbf{X}_{t-1,\mathbf{p}}$ via Eq. (13).
10:            End If
11:      End for
12:      If Posterior sampling then
13:                 Derive the optimized $\mathbf{X}^{\prime}_{t}$ via Eq.  (14) and apply a standard reverse step as Eq. (7) to get the $\mathbf{X}_{t-1}$.
14:      End If
15:End for
16:Return Reconstructed sequence $\mathbf{X}_{0}$.
$\mathcal{I}(\mathbf{p}_{0})=\mathcal{W}\left(\mathbf{X}_{0,\mathbf{p}_{0}},\mathbf{u}_{i}\right),\hskip 1.99997pt\mathbf{u}_{i}=\mathbf{K}\mathbf{D}\Delta\mathbf{p}\mathbf{K}^{-1}\mathbf{u}_{0},$ (9)
where $\mathcal{W}(\cdot,~{}\cdot)$ denotes the image warping function; $\mathbf{u}_{0}$ and $\mathbf{u}_{i}$ are the pixel locations of the views at poses $\mathbf{p}_{0}$ and $\mathbf{p}_{i}$, respectively; $\mathbf{D}$ is the depth map; $\mathbf{K}$ is the camera intrinsic matrix. Moreover, since the ground-truth depth $\mathbf{D}$ is usually not available in practice, we can also use the estimated depth map $\widehat{\mathbf{D}}$ by an off-the-shelf depth estimation method to calculate the estimated pixel location, i.e., $\widehat{\mathbf{u}}_{i}=\mathbf{K}\widehat{\mathbf{D}}\Delta\mathbf{p}\mathbf{K}^{-1}\mathbf{u}_{0}$. By substituting Eq. (9) with an estimated depth map to Eq. (8), we then have
[…]
where $\Delta\mathbf{D}:=\mathbf{D}-\widehat{\mathbf{D}}$, $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$ refers to the warped view from $\mathbf{p}_{0}$ to pose $\mathbf{p}_{i}$ through $\widehat{\mathbf{D}}$, and $\mathcal{E}_{T}$ is a residual term that contains high-order Taylor expansion series and warping error.
Based on the above analysis that the rendered novel view at pose $\mathbf{p}_{i}$ is highly correlated with $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$ and $\bm{\mu}_{t,\mathbf{p}_{i}}$, an effective score function for NVS should take advantage of them. Then, we formulate the score function of NVS-oriented reverse diffusion sampling, which is modulated with the given scene information as
$\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}=\mathop{\mathrm{argmin}}_{\bm{\mu}}\|\bm{\mu}-\bm{\mu}_{t,\mathbf{p}_{i}}\|_{2}^{2}+\lambda(t,\mathbf{p}_{i})\|\bm{\mu}-\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}\|_{2}^{2},$ (11)
where $\lambda(t,\mathbf{p}_{i})>0$ balances the two terms. It is obvious the closed-form solution of Eq. (11) is
$\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}=\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\bm{\mu}_{t,\mathbf{p}_{i}}+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}.$ (12)
With the optimized clean image expectation term $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$, we further utilize two ways to guide the reverse sampling of the pre-trained SVD: (1) Directly Guided Sampling, which is fast with limited quality; and (2) Posterior Sampling, which is slow but more effective. Specifically, the former directly replaces the $\bm{\mu}_{t,\mathbf{p}}$ in Eq. (7) with $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$ as
$d\mathbf{X}_{t,\mathbf{p}_{i}}=\left[\frac{\mathbf{X}_{t,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}}{\sigma(t)}\right]d\sigma(t).$ (13)
While the latter embeds the knowledge of $\widetilde{\bm{\mu}}_{t}$ into $\mathbf{X}_{t}$ via the backward gradient as
$\mathbf{X}^{\prime}_{t}=\mathbf{X}_{t}-\kappa\nabla_{\mathbf{X}_{t}}\|\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t}\|_{2},$ (14)
where $\kappa>0$ controls the updating rate, which is empirically set to $2e^{-2}\sqrt{\sigma(t)}$; $\nabla_{\mathbf{X}_{t}}\|\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t}\|_{2}$ computes the back-propagated gradient on $\mathbf{X}_{t}$; $\mathbf{X}^{\prime}_{t}$ stands for the optimized noised latent $\mathbf{X}_{t}$. Here, we also normalize the gradient $\nabla_{\mathbf{X}_{t}}\|\bm{\mu}_{t}-\widetilde{\bm{\mu}}_{t}\|_{2}$ to stabilize the updating process.
Figure 1: Experimental observations of the relationship (a) between the diffusion estimation error $\mathcal{E}_{D}$ and the noise level $\sigma_{t}$ and (b) between the error of warped image $\mathcal{E}_{T}$ and the changed amount of view pose $\|\Delta\mathbf{p}\|_{2}$.
4.2 Adaptive Determination of $\lambda(t,\mathbf{p}_{i})$
Although we have reformulated the NVS-oriented diffusion process in the preceding section, the value of $\lambda(t,\mathbf{p}_{i})$ in Eqs. (11) and (12), which is crucial to the quality of synthesized views, has to be appropriately determined. Here, we theoretically explore the formulation of $\lambda(t,\mathbf{p}_{i})$ via analyzing the upper boundary of the estimation error of $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$.
Let $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}$ be the ideal value for the score function estimation in the NVS-oriented diffusion process, i.e., the ground-truth view at ${p}_{i}$. Based on the formulation of $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$ in Eq. (12) and the triangle inequality, we have
$\|\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}\leq\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\|\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}.$ (15)
Thus, we propose to minimize the estimation error upper bound to obtain an appropriate and adaptive $\lambda(t,~{}\mathbf{p}_{i})$, i.e.,
$\widehat{\lambda}(t,\mathbf{p}_{i})=\mathop{\mathrm{argmin}}_{\lambda(t,\mathbf{p}_{i})}\frac{1}{1+\lambda(t,\mathbf{p}_{i})}\mathbb{E}_{\mathbf{X}_{t}\sim\mathcal{P}(\mathbf{X}_{t})}(\mathcal{E}_{D})+\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}\mathbb{E}_{\mathbf{X}\sim\mathcal{P}(\mathbf{X})}(\mathcal{E}_{P})+v_{1}|\log(\lambda(t,\mathbf{p}_{i}))|,$ (16)
where $v_{1}>0$ and the last regularization term prevents the weights from being overfitting on the empirically estimated errors. In the following, we will provide the explicit formulations of the two error terms $\mathcal{E}_{D}=\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}$ and $\mathcal{E}_{P}=\|\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}-\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}\|_{2}$  [^2] There is also inherent discretization error as investigated by [25, 26, 60]. However, we could reduce such error terms by enlarging the number of reversing steps., based on theoretical analyses and experimental observations.
Diffusion Estimation Error $\mathcal{E}_{D}$. This error is caused due to the fact that the diffusion model cannot perfectly estimate $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}^{*}$ from the given noised latent $\mathbf{X}_{t}$. Recent works [60, 61] indicate that the derivative $S_{\theta}(\mathbf{X}_{t},t)$ of SDE-based diffusion models varies intensely when $\sigma(t)$ is huge. Accordingly, large values of $\sigma$ would potentially lead to large errors. Moreover, we experimentally investigated the correlation of $\|\mathcal{E}_{D}\|_{2}$ with $\sigma(t)$. As demonstrated in in Fig. 1 (a). $\|\mathcal{E}_{D}\|_{2}$ gradually decreases with the diffusion reverse process (decreasing of $\sigma(t)$). Thus, we empirically formulate $\|\mathcal{E}_{D}\|_{2}=v_{2}\sigma(t)$, where $0<v_{2}$.
Intensity Truncation Error $\mathcal{E}_{P}$. This error is mainly induced by the truncation of the Taylor series (here $\widetilde{\bm{\mu}}^{*}_{t,\mathbf{p}_{i}}$ is the same with $\mathbf{X}$), as shown in Eq. (10). If omitting the high-order terms, we have $\mathcal{E}_{P}\approx\frac{\partial\mathcal{W}\left(\mathbf{X}_{0,\mathbf{p}_{0}},\widehat{\mathbf{u}}_{i}\right)}{\partial\mathbf{u}}\Delta\mathbf{u}+\frac{d\mathcal{I}(\mathbf{p})}{d\mathbf{p}}\Delta\mathbf{p}$. According to the definition of $\Delta\mathbf{u}$ in Eq. (10), we also have $\|\mathcal{E}_{p}\|_{2}\leq v\|\Delta\mathbf{p}\|_{2}$. Together with the experimental observation of the relationship between $\|\mathcal{E}_{p}\|_{2}$ and $\|\Delta\mathbf{p}\|_{2}$ in Fig. 1 (b), we empirically formulate $\|\mathcal{E}_{p}\|_{2}=v_{3}\|\Delta\mathbf{p}\|_{2}$.
Finally, with the explicit formulations of the two error terms, we can rewrite Eq. (16) as
$\widehat{\lambda}(t,\mathbf{p}_{i})=\mathop{\mathrm{argmin}}_{\lambda(t,\mathbf{p}_{i})}\frac{v_{2}\sigma(t)}{1+\lambda(t,\mathbf{p}_{i})}+\frac{\lambda(t,\mathbf{p}_{i})v_{3}\|\Delta\mathbf{p}\|_{2}}{1+\lambda(t,\mathbf{p}_{i})}+v_{1}|\log(\lambda(t,\mathbf{p}_{i}))|,$ (17)
whose closed-form solution is
$\widehat{\lambda}(t,\mathbf{p}_{i})=\frac{-(2v_{1}+Q)+\sqrt{Q^{2}+4v_{1}Q}}{2v_{1}},\quad Q=v_{3}\|\Delta\mathbf{p}\|_{2}-v_{2}\sigma(t).$ (18)
4.3 NVS from Multiviews and Monocular Videos
We have illustrated the single view-based NVS-Solver in Algorithm 1, which can be further modified for NVS from multiple views or monocular videos. Specifically, given multiple views of a static scene and target poses, we first estimate the depth map of each of the given views and derive the warped view at target poses by warping the nearest given views to each target pose. For a monocular video, we warp each frame of the video sequence to the corresponding target pose with the same timestamp. Then, we sample the novel view via the proposed NVS-Solver as Lines 3-16 of Algorithm 1.
5 Experiment
Datasets. For single view-based NVS, we employed totally nine scenes, with six scenes from the Tanks and Temples dataset [22], containing both outdoor and indoor environments. The other three additional scenes are randomly chosen from the Internet. For multiview-based NVS, we used three scenes from the Tanks and Temples dataset [22], including both outdoor and indoor settings, as well as six scenes from the DTU dataset [15], which feature indoor objects. For each scene, we selected two images as input to perform view interpolation. For monocular video-based NVS, we downloaded nine videos from YouTube, each comprising 25 frames and capturing complex scenes in both urban and natural settings.
Implementation Details. We conducted all the experiments with PyTorch using a single NVIDIA GeForce RTX A6000 GPU-48G. We adopted the point-based warping [40] to achieve $\mathcal{W}(\cdot)$ and employed Depth Anything [56] to estimate the maps of the input views. We simultaneously rendered 24 novel views and set the reverse steps as 100 for high-quality sample generation. For the implementation of Eq. (11), since applying directly weighted sum usually results in blurry, we ordered the feature pixels by the $\|\bm{\mu}_{t,\mathbf{p}_{i}}-\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}\|_{2}$ and take the ratio of $\frac{\lambda(t,\mathbf{p}_{i})}{1+\lambda(t,\mathbf{p}_{i})}$ smaller pixels from $\widehat{\mathbf{X}}_{0,\mathbf{p}_{0i}}$ and the others from $\bm{\mu}_{t,\mathbf{p}_{i}}$ to modulate $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}$. We choose the values $(v_{1},v_{2},v_{3})$ as $(1e^{-6},9e^{-1},5e^{-2})$, $(1e^{-6},7e^{-1},1e^{-2})$, and $(1e^{-6},1.75,3e^{-2})$ for single, sparse, dynamic scene view synthesis.
Metrics. We utilized four metrics to measure the reconstruction performance, i.e., Fréchet Inception Distance (FID) [12] evaluating the quality and diversity of synthesized views; Absolute trajectory error (ATE) [10] measuring the difference between the estimated trajectory of a camera or robot and the ground truth trajectory; Relative pose error (RPE) [10] measuring the drift, where we separately calculated the transition and rotation as RPE-T and RPE-R, respectively. Due to that the current depth estimation algorithms are hard to derive the absolute depth, we refrain from calculating image metric values like PSNR or SSIM.
Table 1: Quantitative comparison of different methods on view synthesis, where we measure the FID and the error of view pose. We name our method with NVS-Solver (DGS) or (Post) for utilizing directly guided sampling in Eq. (13) or posterior sampling in Eq. (14), respectively. "*" indicates incomplete evaluation, where the corresponding metric cannot output a meaningful value on some sequences by the method due to the failure of pose estimation. “– –” denotes that the method cannot work on the condition or the metrics cannot be calculated. For all metrics, the lower, the better.
Methods Train Single view Multi-view
FID ATE RPE-T RPE-R FID ATE RPE-T RPE-R
Sparse Gaussian [55] 369.19 – – – – – – 324.60 – – – – – –
Sparse Nerf [49] – – – – – – – – 180.26 6.129 1.711 1.804
Text2Nerf [59] 187.05 2.223 0.718 0.107 – – – – – – – –
3D-aware [54] $\times$ 217.19 2.836 1.258 1.662 211.25 2.159 5.679 2.119
MotionCtrl [51] $\times$ 179.24 3.851 0.705 0.835 154.27 37.68 19.61 1.646
NVS-Solver (DGS) $\times$ 166.50 4.533 0.810 0.742 124.31 22.00 17.34 1.338
NVS-Solver (Post) $\times$ 165.12 0.767 0.156 0.170 126.44 4.052 2.030 0.330
Figure 2: Visual comparison of single view-based NVS results by (a) Text2Nerf [59], (b) 3D-aware [54], (c) MotionCtrl [51], (d) our NVS-Solver (Post). The middle view of each scene highlighted with the red rectangle refers to the input view. Here, we only show the results of the best two of all compared methods. We also refer reviewers to the Appendix and video demo contained in the supplementary file for more impressive results and comparisons.
Figure 3: (a) The two input views of each scene highlighted with the red rectangle. Visual results of multiview-based NVS by (b) 3D-aware [54], (c) MotionCtrl [51], (d) Our NVS-Solver (Post).
5.1 Results of NVS from Single or Multiple Views of Static Scenes
Fig. 2 visualizes synthesized novel views of two scenes by different methods from single views, where we can see that our NVS-Solver can consistently generate high-quality novel views with visually pleasing geometry and textures. The quantitative results listed in Table 1 also validate the significant superiority of our method over state-of-the-art methods. Although MontionCtrl [51] can generate high-quality images reflected by the FID value, its results have significantly large view pose errors reflected by the higher ATE and RPE values. For Text2Nerf [59], only 7 out of the total 9 sequences can be calculated metrics, which may be induced by the inherent errors of the learned geometric structure.
Fig. 3 shows the synthesized views by different methods from two given views, where it can be seen that our NVS-Solver outperforms state-of-the-art methods by clearer views, especially for the $2^{nd}$ and $3^{rd}$ scenes. In addition, it is worth noting that the lower ATE of 3D-aware [54] is due to the incomplete evaluation (See the supplementary material for more explanations).
Figure 4: Visual comparison on dynamic scene view synthesis of (a) input frames in the corresponding time of generated images, (b) Deformable-Gaussian [57], (c) 4D-Gaussian [52], (d) 3D-aware [54], (e) MotionCtrl [51], (f) Our NVS-Solver (Post).
5.2 Results of NVS from Monocular Videos of Dynamic Scenes
Table 2: Quantitative comparison of different methods on NVS from monocular videos of dynamic scenes. For all metrics, the lower, the better.
Methods Train FID ATE RPE-T RPE-R
Deformable-Gaussian [57] 115.82 1.813 0.678 0.613
4D-Gaussian [52] 74.34 2.087 0.625 0.825
3D-aware [54] $\times$ 159.03 3.100 1.343 1.368
MotionCtrl [51] $\times$ 70.35 3.384 1.069 0.653
NVS-Solver (DGS) $\times$ 37.973 2.236 0.691 0.446
NVS-Solver (Post) $\times$ 39.86 2.308 0.725 0.400
The non-generative Gaussian-based methods Deformable-Gaussian [57] and 4D-Gaussian [52] usually cannot handle the marginal area, as illustrated in $2^{nd}$, $4^{th}$ and $6^{th}$ columns of Fig. 4 (b) and (c). Although the SVD-based method MotionCtrl can generate the boundary region as shown in Fig. 4 (e), the synthesized views are blurry and the poses of generated samples cannot follow the prompt, especially on the $3^{rd}$ sample. However, our method consistently generates high-quality novel views with lower pose errors, indicating its strong potential.
5.3 Ablation Study
Table 3: Quantitative comparison of different numbers of inference steps. All metrics are the lower the better.
Inference step FID ATE RPE-T RPE-R
25 175.432 5.317 0.691 0.847
50 168.938 2.275 0.428 0.402
100 165.12 0.767 0.156 0.170
Reverse inference steps. The quantitative results in Table 3 show that the synthesized image quality of our method does not improve intensely with the number of inference steps increasing. However, the pose error decreases significantly, indicating the necessity of a sufficient number of inference steps for accurately rendering novel views.
Posterior sampling vs. Directly guided sampling. The quantitative results listed in Tables 1 and 2 show that both the NVS-Solver (DGS) and NVS-Solver (Post) can generate high-quality images with comparable FID. However, the view pose of NVS-Solver (Post) is much more accurate than NVS-Solver (DGS), which is also visually verified by the results in Figs. 5 (c) and (d). Besides, NVS-Solver (DGS) takes 6 minutes for rendering 25 views, while NVS-Solver (Post) uses 1 hour.
Figure 5: Ablation studies of the proposed weight strategy of $\widehat{\lambda}(t,\mathbf{p}_{i})$ and embedding strategy. Visual comparison of (a) warped input image $\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$, (b) results without our proposed weight strategy, (c) results of our NVS-Solver (DGS), and (d) results of our NVS-Solver (Post).
Figure 6: NVS results of our NVS-Solver (Post) with different trajectories, where the given view is bounded by a blue circle. We draw the pose at the right bottom corner of each sub-figure.
Effectiveness of adaptive $\widehat{\lambda}(t,\mathbf{p}_{i})$. Here, we illustrate the effectiveness of adaptively adjusting $\widehat{\lambda}(t,\mathbf{p}_{i})$ via setting it to +$\infty$+ ∞, i.e., $\widetilde{\bm{\mu}}_{t,\mathbf{p}_{i}}=\widehat{\mathbf{X}}_{0,\mathbf{p}_{i}}$. Visual comparison results in Fig. 5 indicates that the proposed method significantly correct the warping errors ($1^{st}$, $2^{nd}$ and $4^{th}$ samples in Fig. 5) and non-Lambert reflection ($3^{rd}$ sample in Fig. 5) as indicated in $\mathcal{E}_{T}$ of Eq. (10).
Different Trajectories. We detailedly visualize the result of the proposed method on different trajectories. We apply transitions on eight different directions. The results are shown in Fig. 6. The proposed method can consistently render high-quality novel views.
6 Conclusion & Discussion
We have presented a training-free novel view synthesis paradigm, namely NVS-solver. NVS-Solver achieves remarkable performance compared with state-of-the-art methods. The advantages are credited to the powerful generative capacity of the pre-trained large stable video diffusion model and our elegant designs of the adaptive modulation of the diffusion score function through comprehensive theoretical and empirical analyses.
Although our NVS-Solver takes longer than existing methods, the promising generative capacity inherent in the large pre-trained diffusion model may attract improvements in the future. Since the proposed method can synthesize views more accurate poses, we believe it may be a potential solution for the distillation of a pose controllable video diffusion model. In the age of generative intelligence, the authors expect the proposed algorithm would inspire future work to unify computer graphics pipelines with generative models. Finally, the authors sincerely appreciate Stability AI for the open-sourced SVD [5].
Appendix A Visual Comparison Details
In this section, we visually demonstrate the results as shown in Figs. A-1, A-2, and A-3. The visual result demonstrates the superiority of the proposed method.
Figure A-1: Visual comparison of single view-based NVS results by (a) Text2Nerf [59], (b) 3D-aware [54], (c) MotionCtrl [51], (d) our NVS-Solver (Post). The middle view of each scene highlighted with the red rectangle refers to the input view.
Figure A-2: (a) The two input views of each scene highlighted with the red rectangle. Visual results of multiview-based NVS by (b) 3D-aware [54], (c) MotionCtrl [51], (d) Our NVS-Solver (Post).
Figure A-3: Visual comparison on dynamic scene view synthesis of (a) input frames in the corresponding time of generated images, (b) Deformable-Gaussian [57], (c) 4D-Gaussian [52], (d) 3D-aware [54], (e) MotionCtrl [51], (f) Our NVS-Solver (Post).
Appendix B Quantitative Comparison Details
We give the detailed quantitative comparisons of different methods as shown in Tables B-1, B-2 and B-3. The results demonstrate that the proposed methods, NVS-Solver (DGS) and (Post) outperform SOTA methods in most scenes.
Table B-1: Quantitative comparison of different methods on single view synthesis. For all metrics, the lower, the better.
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl NVS-Solver (DGS) NVS-Solver (Post)
Auditorium ATE – – – – 4.424 2.850 7.111 17.160 1.261
RPE-T – – – – 2.020 3.398 1.448 2.669 0.265
RPE-R – – – – 0.266 2.609 1.320 1.192 0.144
Barn ATE – – – – 1.133 2.832 3.217 2.638 0.430
RPE-T – – – – 0.307 0.868 0.369 0.533 0.104
RPE-R – – – – 0.064 2.336 1.074 0.632 0.197
Castle ATE – – – – – – 2.831 3.045 3.038 0.883
RPE-T – – – – – – 1.425 0.716 0.634 0.171
RPE-R – – – – – – 1.036 0.483 0.983 0.039
Church ATE – – – – 2.167 2.830 3.256 3.017 0.664
RPE-T – – – – 0.486 0.860 0.798 0.636 0.107
RPE-R – – – – 0.093 2.476 1.553 1.000 0.176
Family ATE – – – – 1.264 2.830 1.941 0.944 1.023
RPE-T – – – – 0.377 0.867 0.378 0.212 0.191
RPE-R – – – – 0.050 2.295 0.890 0.461 0.514
Ignatius ATE – – – – 1.561 2.831 3.431 1.226 0.408
RPE-T – – – – 0.365 0.866 0.495 0.342 0.129
RPE-R – – – – 0.050 2.298 1.043 0.894 0.243
Palace ATE – – – – – – 2.835 3.903 0.959 1.062
RPE-T – – – – – – 0.904 0.517 0.221 0.192
RPE-R – – – – – – 0.578 0.363 0.313 0.101
Seaside ATE – – – – 2.385 2.830 4.007 5.292 0.395
RPE-T – – – – 0.754 0.717 0.522 0.639 0.090
RPE-R – – – – 0.073 0.473 0.283 0.311 0.040
Trees ATE – – – – 2.628 2.854 4.740 6.521 0.776
RPE-T – – – – 0.718 1.420 1.098 1.399 0.155
RPE-R – – – – 0.151 0.859 0.506 0.894 0.079
Table B-2: Quantitative comparison of different methods on sparse view synthesis. For all metrics, the lower, the better.
Scene Metric Sparse Gaussian Sparse Nerf Text2Nerf 3D-aware MotionCtrl NVS-Solver (DGS) NVS-Solver (Post)
caterpillar ATE – – 1.697 – – – – 2.826 2.250 0.330
RPE-T – – 0.046 – – – – 0.040 0.010 0.002
RPE-R – – 2.337 – – – – 0.680 0.305 0.032
playground ATE – – 2.813 – – 0.721 1.918 0.597 0.033
RPE-T – – 0.058 – – 0.059 0.029 0.008 0.001
RPE-R – – 0.156 – – 3.534 0.524 0.201 0.113
truck ATE – – 2.556 – – – – 2.718 0.058 0.068
RPE-T – – 0.142 – – – – 0.088 0.010 0.002
RPE-R – – 11.96 – – – – 0.954 0.287 0.048
scan1 ATE – – 7.980 – – 2.872 – – 33.29 7.069
RPE-T – – 2.518 – – 8.477 – – 31.02 3.918
RPE-R – – 0.302 – – 1.411 – – 2.995 0.592
scan2 ATE – – 8.318 – – – – 48.00 6.556 4.480
RPE-T – – 2.609 – – – – 31.60 5.902 2.211
RPE-R – – 0.307 – – – – 1.614 0.648 0.324
scan3 ATE – – 8.795 – – – – 46.22 8.660 7.617
RPE-T – – 2.633 – – – – 31.20 6.104 3.976
RPE-R – – 0.269 – – – – 1.656 0.940 0.564
scan5 ATE – – 7.902 – – 2.886 53.38 59.03 3.436
RPE-T – – 2.510 – – 8.495 16.13 32.65 2.270
RPE-R – – 0.297 – – 1.413 1.526 2.551 0.326
scan15 ATE 7.072 – – – – 48.75 55.68 6.504
RPE-T 2.394 – – – – 30.57 42.98 2.796
RPE-R 0.327 – – – – 1.692 1.690 0.352
scan55 ATE 8.025 – – – – 97.60 31.86 6.936
RPE-T 2.487 – – – – 47.21 37.41 3.098
RPE-R 0.275 – – – – 4.520 2.422 0.618
Table B-3: Quantitative comparison of different methods on dynamic view synthesis. ’_1/_2’ mean the we synthesize novel views by moving the camera on the right/left circle trajectory. For all metrics, the lower, the better.
Scene Metric Deformable-Gaussian 4D-Gaussian 3D-aware MotionCtrl NVS-Solver (DGS) NVS-Solver (Post)
Street_1 ATE 2.834 2.817 2.847 2.746 0.619 0.619
RPE-T 1.463 1.342 1.383 0.811 0.306 0.307
RPE-R 1.009 1.322 1.0340 0.906 0.217 0.218
Street_2 ATE 0.486 0.819 1.470 – – 2.842 2.842
RPE-T 0.169 0.163 0.235 – – 0.647 0.647
RPE-R 0.439 0.800 0.528 – – 0.604 0.602
Kangaroo_1 ATE – – 3.546 4.193 1.396 3.421 3.131
RPE-T – – 0.979 1.347 0.778 1.501 1.482
RPE-R – – 0.361 2.637 0.606 0.574 0.444
Kangaroo_2 ATE 3.302 2.438 4.099 5.681 5.922 4.248
RPE-T 1.672 0.852 2.489 1.973 2.065 1.608
RPE-R 0.639 0.372 0.954 0.303 0.791 0.353
Train3_1 ATE 1.658 1.187 4.667 5.626 3.744 1.130
RPE-T 0.698 0.650 2.642 1.501 0.999 0.402
RPE-R 0.468 0.588 1.494 0.471 0.455 0.245
Train3_2 ATE 1.144 2.880 4.031 2.186 1.026 1.315
RPE-T 0.968 1.108 1.748 0.908 0.448 0.476
RPE-R 0.785 1.483 2.266 0.574 0.567 0.529
Train5_1 ATE 2.838 1.600 3.011 3.274 0.624 0.815
RPE-T 0.313 0.252 1.537 0.880 0.196 0.165
RPE-R 0.546 0.699 1.523 1.042 0.453 0.629
Train5_2 ATE 1.046 2.3516 2.708 3.3222 0.851 2.874
RPE-T 0.335 0.377 0.513 0.719 0.1916 0.991
RPE-R 0.734 1.686 1.347 0.666 0.599 0.653
Road_1 ATE 0.521 0.548 1.195 2.783 0.774 3.328
RPE-T 0.219 0.251 0.667 1.061 0.326 1.016
RPE-R 0.330 0.363 0.679 0.479 0.100 0.239
Road_2 ATE 2.492 2.679 2.775 3.445 2.534 2.781
RPE-T 0.260 0.275 0.867 0.991 0.230 0.158
RPE-R 0.564 0.579 1.214 0.828 0.099 0.091